suppose a businessman has to rank (in order) different locations to start a new business. If he has a total of 12 locations to choose from, how many ways can he rank 4 of those locations?
Answer by stanbon(75887)

If he has a total of 12 locations to choose from, how many ways can he rank 4 of those locations?
---
Rank implies order.
Order implies permutations.
---
Your problem:
12P4 = 12!/(12-4)! = 12*11*10*9 = 11,880
============
Cheers,
Stan H.
============